OpenStudy (anonymous):

Pls help me solve for all real values of x only.. 3x^4 + 3x^3 - 21x^2 - 3x + 18 = 0

OpenStudy (anonymous):

manual way: change of variable, reduce poly auto: wolfram, solve for x

OpenStudy (anonymous):

thanks..

OpenStudy (precal):

synthetic division also works

OpenStudy (anonymous):

thanks again

OpenStudy (precal):

4th power means you are looking for 4 solutions

OpenStudy (anonymous):

yeap, syntheic rules, is like factoring polys

OpenStudy (anonymous):

yup. thanks, guys! :)

OpenStudy (anonymous):

The synthetic division method factors it fast.

OpenStudy (anonymous):

\[0=3x^4 + 3x^3 - 21x^2 - 3x + 18=(x-1)(x-2)(x+3)(x+9)\]

OpenStudy (anonymous):

So we end up with solutions of \[x \in \left\{ -9,-3,1,2 \right\}\]

OpenStudy (anonymous):

thank you! :D